OpenStudy (anonymous):

a bullet going with speed 350m/s enters a concrete wall and penerate a distance of 0.5cm before coming to rest.find decelaration.

OpenStudy (anonymous):

You have initial velocity, final velocity, and displacement. Therefore, you could use: \[a=\frac{v_1-v_0}{t}\]and\[s=v_0t+\frac{1}{2}at^2\], which should allow you to find t, which would further allow you to find a!

OpenStudy (anonymous):

\[v ^{2} - u ^{2}= 2 a s\]

OpenStudy (anonymous):

- 350 * 350 = 2 * 1/ 200 * a Or, a = - 350 * 350 * 100

OpenStudy (anonymous):

@Taufique bhai, please help.

OpenStudy (anonymous):